How many liters of nitric acid solution, with a concentration of 0.05 mol / dm3, can be made if we have 150 cm3 of 60% nitric acid, whose density is 1.37 g / cm3?


1
Expert's answer
2021-05-14T02:34:47-0400

Solution:

Calculate the mass of nitric acid solution:

Mass of HNO3 solution = Solution volume × Density of solution

Mass of HNO3 solution = 150 cm3 × 1.37 g/cm3 = 205.5 g


Calculate the mass of nitric acid (HNO3) in the solution:

Mass of HNO3 = Mass of HNO3 solution × w(HNO3)

Mass of HNO3 = 205.5 g × 0.6 = 123.3 g


Calculate the moles of nitric acid (HNO3) in the solution:

Moles of HNO3 = Mass of HNO3 / Molar mass of HNO3

The molar mass of HNO3 is 63.01 g mol-1.

Moles of HNO3 = (123.3 g) / (63.01 g mol-1) = 1.9568 mol


Calculate liters of nitric acid solution:

Liters of HNO3 solution = Moles of HNO3 / Molarity of HNO3

Liters of HNO3 solution = (1.9568 mol) / (0.05 mol dm-3) = 39.137 dm3 = 39.14 L


Answer: 39.14 L of nitric acid solution.